OpenStudy (anonymous):

solve: 3x^2+5x=0 (i need to know how to do plz, not just answer...)

OpenStudy (anonymous):

factor out the common factor of x as a start \[3x^2+5x=0\] \[x(3x-5)=0\]

OpenStudy (anonymous):

then set each factor = 0 and solve. \[x=0\] or \[3x+5=0\] \[3x=-5\] \[x=-\frac{5}{3}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw